OpenStudy (lstrasmann):

Find the volume of a sphere with diameter 22 meters. A. 6272.19 m3 B. 5575.28m3 C. 2027.37 m3 D. 44602.24 m3

OpenStudy (raden):

Volume of a sphere = 4/3 * pi * r^3 with r is the radius

OpenStudy (lstrasmann):

OK, thanks a lot.. I think I understand now.

OpenStudy (raden):

u have to convert the diameter to be the radius, it is r = 1/2 diameter

OpenStudy (lstrasmann):

So it would be\[\frac{ 4 }{ 3 }*\pi*11^3 = 5575.2797\]Which is B

OpenStudy (lstrasmann):

Thanks for helping

OpenStudy (raden):

yes, that's right you're welcome ;)